OpenStudy (anonymous):

The edges of two similar pyramids have a scale factor of 2 : 5. If the volume of the smaller pyramid is 16, what is the volume of the larger pyramid?

OpenStudy (anonymous):

sot the answer is 40 because i don't have that answer choice

OpenStudy (mayankdevnani):

yaa @mark_o. you are right..

OpenStudy (anonymous):

are you ever right @mayankdevnani ?

OpenStudy (mayankdevnani):

no... but @mark_o. was right

OpenStudy (mayankdevnani):

got it @Algebraic!

OpenStudy (anonymous):

is ans 250

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

simple... vol. of similar pyramids is proportional to the cubes of the sides....

OpenStudy (anonymous):

im sorry that was cube VL/Vs=(5/2)^3 VL=(5/2)^3Vs VL=(5/2)^3 (16) VL=250 ...